OpenStudy (anonymous):

how do I complete the square for 2x^2+9x+4=0

OpenStudy (cydney_morgan):

The expression you have given is not an equation (does not equal something), so it cannot be solved. I am going to assume that the expression is supposed to equal zero, and then solve that equation. The equationthat I am going to solve is 2x2 - 9x + 4 = 0. The method I am going to use is called completing the square. 2x2 - 9x + 4 = 0 2x2 - 9x + 4 -4 = 0 -4 Subtract 4 from each side 2x2 - 9x = -4 Simplify (2x2 - 9x)/x = -4/2 Divide both sides by two (the coefficient of x2) x2 -9/2*x = -2 Distribute on the left, simplify the right x2 - 9/2*x + 81/16 = -2 + 81/16 Add 81/16 to each side (divide the coefficient on x by two, then square the result) (x - 9/4)2 = -2 + 81/16 Factor the left side as a perfect square (9/4 is the square root of 81/16) (x - 9/4)2 = -32/16 + 81/16 Change the fractions on the right so that they have common denominators (x - 9/4)2 = 49/16 Simplify x - 9/4 = +/- sqrt(49/16) Take the square root of both sides (sqrt means square root) x - 9/4 = +/- 7/4 Simplify x - 9/4 = 7/4 or x -9/4 = -7/4 Separate the problem into positive and negative values on the right x - 9/4 + 9/4 = 7/4 +9/4 or Add 9/4 to both sides on both possible solutions x - 9/4 + 9/4 = -7/4 + 9/4 x = 16/4 or x = 2/4 Simplify each solution x = 4 or x = 1/2 Reduce the fractions Completing the square is a valuable tool to use, because any binomial equation can be solved with it. Even though you could find the answer to this problem without it, I wanted to give you a chance to learn how to use it.

OpenStudy (anonymous):

The question is asking me to complete the square for 2x^2+9x+4=0; help me please! this is my first time using service.

OpenStudy (wolf1728):

2x² + 9x +4 =0 To complete the square: Move the "non X" term to the right: 2x² + 9x =-4 Divide the equation by the coefficient of X² x² + 4.5x = -2 take the coefficient of X and divide it by 2 then square it 4.5 / 2 = 2.25 and 2.25² = 5.0625 add 5.0625 to both sides of the equation x² + 4.5x +5.0625= -2 +5.0625 Take square root of both sides x + 2.25 = sqrt(3.0625) x + 2.25 = 1.75 x = -.5 sqrt of 3.0625 also equals -1.75 x + 2.25 = -1.75 x = -4